1. Using excel find the answers to the following: (remember to use excel functions)

a. Find the price today of a five year bond with a $1000 face value with a 10 percent coupon. Assume that investors require an 8 percent YTM. Assume also that payments are made semi-annually.

b. Suppose the bond in problem 1 sells for $1200 today. Solve for its YTM.

c. Find the price of a share of common stock that is expected to pay a dividend of $2 next year. Dividends are expected to grow at 4 percent through year 3, then at 3 percent for 2 years, and then 2 percent forever. Assume that stockholders require a 10 percent return.                    
Answer #1

1.
=PV(8%/2,2*5,-10%*1000/2,-1000)=1081.10895779355

2.
=RATE(2*5,10%*1000/2,-1200,1000)*2=5.3841694502009%

3.
=2/1.1+2*1.04/1.1^2+2*1.04^2/1.1^3+2*1.04^2*1.03/1.1^4+2*1.04^2*1.03^2/1.1^5+2*1.04^2*1.03^2/1.1^5*1.02/(10%-2%)=26.2776804863056